OpenStudy (anonymous):

write an equation for y=3/x that has the asymptotes x=-8 y=6

OpenStudy (anonymous):

@coolsday @Christythecreatorr

OpenStudy (anonymous):

What do you think it is?

OpenStudy (anonymous):

the equation y=1/x has horizontal and vertical asymptotes at 0

OpenStudy (anonymous):

If we translate the equation y=1/x up by 6 units, the vertical asymptote will still be 0. However, the horizontal asymptote will change to 6. The horizontal asymptote approaches the value y=6, but will never touch it.

OpenStudy (anonymous):

is it D...? y= 3/(x+8) -6

OpenStudy (anonymous):

It is not D, because your horizontal asymptote is y=6. Therefore, if it was D, your graph would have a horizontal asymptote at y=-6

OpenStudy (anonymous):

so A?

OpenStudy (anonymous):

i mean C...

OpenStudy (anonymous):

Why is it C?

OpenStudy (anonymous):

no b? cuz its +6 and -8?

OpenStudy (anonymous):

yeah, so the equation is C